# Explain Like I’m Five: Using LLMs to Improve PDE Surrogate Models with Text

### v4

@@ Abstract
Solving Partial Differential Equations (PDEs) is ubiquitous in science and engineering. Computational complexity and difficulty in writing numerical solvers has motivated the development of machine learning techniques to generate solutions quickly. Many existing methods are purely data driven, relying solely on numerical solution fields, rather than known system information such as boundary conditions and governing equations. However, the recent rise in popularity of Large Language Models (LLMs) has enabled easy integration of text in multimodal machine learning models. In this work, we use pretrained LLMs to integrate various amounts known system information into PDE learning. Our multimodal approach significantly outperforms our baseline model, FactFormer, in both next-step prediction and autoregressive rollout performance on the 2D Heat, Burgers, Navier-Stokes, and Shallow Water equations. Further analysis shows that pretrained LLMs provide highly structured latent space that is consistent with the amount of system information provided through text.
@@ 1 Introduction
Recently, machine learning methods have exploded in popularity to address these downsides in traditional numerical solvers. The primary aim is to reduce time-to-solution and bypass expensive calculations. Physics Informed Neural Networks (PINNs)Raissi et al. (2019) have been widely successful in small-scale systems. More recently, neural operators(Li et al., 2021; Lu et al., 2021; Li et al., 2023a) have improved upon PINNs, showing promise for larger scale, general purpose surrogate models. However, these models generally are purely data-driven, and do not use any additional, known, system information.
Additionally, owing to the success of large language models (LLMs) in many other domains, such as robotics(Kapoor et al., 2024; Bartsch & Farimani, 2024), design(Kumar et al., 2023; Badagabettu et al., 2024; Jadhav & Farimani, 2024; Jadhav et al., 2024), some works have begun incorporating LLMs into the PDE surrogate model training pipeline. Universal Physics Solver (UPS)(Shen et al., 2024) incorporates pretrained LLMs, but uses limited text descriptions that do not fully utilize LLM capabilities. Unisolver(Zhou et al., 2024b) takes in a LaTeX description of the system as a prompt, but uses MLP encoders for additional system information such as boundary conditions, which does not fully explore the capabilities of the LLM. ICON-LM(Yang et al., 2024) uses longer text descriptions, but trains the LLM to make numerical predictions from input data and captions, adding additional complexity to the model architecture. Additionally, by using softmax-based attention for numerical predictions, benchmarks are limited to 1D.
Contributions: In this work, we aim to more fully utilize LLM understanding in PDE surrogate modeling and incorporate text information into neural operators. To that end, we introduce a novel multimodal PDE framework given in figure 1. This framework is built on top of a FactFormer(Li et al., 2023b) and successfully incorporates system information through text descriptions. The cross-attention based multimodal block is given in figure 2. We benchmark this framework on the Heat, Burgers, Navier-Stokes, and Shallow Water equations to provide a wide variety of physical behavior. Various boundary conditions, initial conditions, and operator coefficients are used, making these benchmarks more challenging than existing data sets. We test various levels of system information, and incorporate this as conditioning information into FactFormer.
@@ 2 Related Work
Neural Solvers: Neural solvers have been developed largely to counter the drawbacks of traditional numerical methods. Physics Informed Neural Networks (PINNs)(Raissi et al., 2019) incorporate governing equations through a soft constraint on the loss function. PINNs have been shown to be effective in small-scale problems, but tend to be difficult to train(Wang et al., 2022; Rathore et al., 2024). Neural operators(Kovachki et al., 2023) were developed and show improvement over PINNs on a large variety of PDE learning tasks. Based on the universal operator approximation theorem, neural operators learn a functional that maps input functions to solution functions. Neural operators such as Fourier Neural Operator (FNO)(Li et al., 2021), based on kernel integral blocks in Fourier space, DeepONetLu et al. (2021), based on parameterizing the input functions and an embedding of the points at which the functions are evaluated, and OFormer(Li et al., 2023a), based on softmax-free attention. Despite very promising results, neural operators are usually purely data-driven and do not use any system information outside of solution fields.
Utilizing System Information While these neural operators perform well, they are often purely data driven. More recent works have incorporated various different aspects of the governing systems. Takamoto et al. (2023a) incorporates operator coefficients(Takamoto et al., 2023a) through the CAPE module. While coefficients are important in determining system behavior, system parameters such as boundary conditions, forcing terms, and geometry can play an equally important role, and cannot be incorporated through CAPE. Further work incorporates governing equations into models. Lorsung et al. (2024) developed the PITT framework that uses a transformer-based architecture to process system information as text. Additionally, PROSE(Liu et al., 2023) poses a multi-objective task to both make a prediction and complete partially correct governing equation. However, these frameworks rely on notational consistency between samples, and do not offer an easy integration of different geometries. Lastly, Hao et al. (2023) introduced novel Heterogeneous Normalized Attention and Geometric Gating mechanisms for flexible GNOT model. GNOT is able to incorporate different system information, such as coefficients and geometry. This flexible architecture can incorporate many different system parameters, but requires additional implementation details for each additional modality, and may not be able to capture qualitative aspects of systems, such as flow regime.
@@ 3 Data Generation
We generate data from the Heat, Burgers, and Incompressible Navier Stokes equations, which are popular benchmarks for fluid dynamics surrogate models. Additionally, we benchmark on the Shallow Water equations from PDEBench(Takamoto et al., 2023b). Data for the Heat and Burgers equations are generated using Py-PDE(Zwicker, 2020) due to ease of simulating different boundary conditions, and data is generated for Navier Stokes using code from Fourier Neural Operator(Li et al., 2021). A diverse data set is generated in order to create a more challenging setup than existing benchmarks, that often have the same operator coefficients and boundary conditions for all samples, only varying the initial conditions. In our case, we use different initial conditions, operator coefficients, and boundary conditions. While existing data sets offer distinct challenges with regards to governing equations, they do not offer a lot of data diversity with regard to system parameters, which LLMs are well-suited to handle. PDEBenchTakamoto et al. (2023b), for example, has multiple 2D data sets that represent a variety of physical processes. However, only eight different operator coefficient combinations are used for the 2D Compressible Navier Stokes equations, all with the same boundary conditions. In order to fully utilize the capabilities of pretrained LLMs, as well as present a more challenging benchmark, we vary boundary conditions and operator coefficients for the Heat and Burgers equations in 2D, and vary viscosity and forcing term amplitude for the Incompressible Navier-Stokes equations in 2D.
@@ 3.1 Heat Equation
We generate data on a simulation cell given by: on a 64x64 grid with boundary conditions sampled from . The values for the Neumann and Dirichlet boundary conditions are sampled uniformly: . All four walls have the same boundary condition type and value for a given simulation. Initial conditions are chosen from three different distributions: exponential field given by , sum of sine and cosine given by , and product of sines given by for . Lastly, our diffusion coefficient was sampled randomly according to . We simulated each trajectory for 2 seconds on a 64x64 grid. These distribution of diffusion coefficients was chosen so the diffusive dynamics behaved on approximately the same scale as the advection dynamics of Burgers Equation given below.
@@ 3.2 Burgers’ Equation
Burgers equation models shock formation in fluid waves, given below in equation 2, where we are predicting the [height], at each time step.
@@ 3.3 Incompressible Navier-Stokes Equations
The Navier-Stokes data generated here follows the setup from Lorsung et al. (2024). Our simulation cell is given by: with periodic boundary conditions. Our viscosity is sampled according to , and our forcing term amplitude is sampled according to . The initial condition is given by a Gaussian random field. Data generation is done on a 256x256 grid that is evenly downsampled to a 64x64 grid for this work.
@@ 3.4 Shallow-Water Equations
Lastly, we take the Shallow-Water data set from PDEBench(Takamoto et al., 2023b). In this case, we are predicting the water height at each time.
@@
For each combination of initial conditions and boundary conditions, 900 samples were generated with randomly generated coefficients. For Navier Stokes, we did the vorticity formulation, from FNO, and varied viscosity and forcing term amplitude, with random gaussian field initial condition. 30 second simulation. Also used PDEBench’s shallow water data set. These are all single-channel. In our experiments, B refers to incorporating boundary condition information through text, C refers to incorporating coefficient information through text, and Q refers to incorporating qualitative information through text.
@@ 4.1 System Descriptions
Burgers equation models a conservative system that can develop shock wave discontinuities. Burgers equation is a first order quasilinear hyperbolic partial differential equation.
All of our basic system descriptions come from Wikipedia entries for their respective equations(Wikipedia, 2024a; b; c; d). Note: Burgers’ system is first order quasiliner hyperbolic only in the inviscid case, where we are working in the viscous case. The second sentence should instead read: Burgers equation is a second order partial differential equation.Additional testing has shown in appendix G that this misnomer does not affect results. Next, boundary condition information is added, in this case Neumann boundary conditions:
@@ 4.2 Multimodal PDE Learning
The backbone of our multimodal surrogate model is the FactFormer(Li et al., 2023b) and our framework is given in figure 1. FactFormer was chosen because it provides a fast and accurate benchmark model. We add our system description as conditioning information both before and after factorized attention blocks. FactFormer is a neural operator that learns a functional that maps from input function space to solution function space as , with parameters . In our case, we are specifically learning the functional conditioned on our system information . For a given input function evaluated at points , we are leaning the the operator given in equation 5. That is, our network learns to make predictions for our solution function also evaluated at points .
@@
This multimodal block is agnostic to both the data-driven backbone as well as the LLM. We compare Llama 3.1 8B(AI@Meta, 2024) and all-mpnet-base-v2 from the Sentence-Transformer package(Reimers & Gurevych, 2019) as our pretrained LLMs. Llama 3.1 8B was used because of its good performance across a wide variety of benchmarks, as well as its small size allowing us to use only a single GPU to generate sentence embeddings. The word embeddings are averaged for each sentence to provide a single embedding that can be used in our projection layer. all-mpnet-base-v2 was used because it is designed to generate embeddings from sentences that are useful for tasks like sentence similarity. In all of our experiments, the LLM is frozen and a projection head is trained. This significantly improves training time by allowing us to generate the sentence embeddings before training, and avoiding expensive gradient computations for our LLMs.
@@ 5 Results
We benchmark our multimodal model against its baseline variant on a number of challenging tasks. Our data vary the distribution of initial conditions, operator coefficients, and boundary conditions, which provides a much more challenging setting than many existing benchmarks. In each experiment, the combined data set is the Heat, Burgers, and Navier-Stokes data sets, where Shallow Water is used solely for fine-tuning, with a 90-10-10 train-validation-test split. 4,000 samples per equation are used for Heat, Burgers, and Navier-Stokes data unless otherwise noted, and we use the entire 1,000 sample Shallow-Water data set when using more samples for our other data sets. We use one frame of data to predict the next frame, which allows us to asses how well our multimodal approach captures known system information, which is the operator . That is the input function at points and time conditioned on sentence information is used to predict our solution function at time . Using only a single step of input here presents the additional challenge that it is difficult to infer boundary conditions and impossible to infer operator coefficients from only a single frame of data. Our temporal horizon is limited to 21 steps for each equation during training (initial condition and 20 steps), and 40 steps during autoregressive rollout evaluation. Relative Error(Li et al., 2021) is used for both training and next-step evaluation in all of our experiments. Autoregressive rollout error is reported as Mean Squared Error. In our results, we used SentenceTransformer (ST)(Reimers & Gurevych, 2019) and Llama 3.1 8B (Llama)(AI@Meta, 2024), using various combinations of boundary condition (B), coefficient (C) and qualitative (Q) information. Each reported result is the mean and standard deviation across three random seeds.
@@ 5.1 Next-Step Prediction
First, we evaluate next-step predictive accuracy, where we take one frame of data and use it to predict the next across our entire temporal window. We see in figure 3 that our multimodal FactFormer significantly outperforms baseline FactFormer when using SentenceTransformer for our Heat, Burgers, and Shallow Water data sets, both with and without transfer learning, where we have small improvement on our Navier-Stokes benchmark when doing standard training. Transfer learning was done because Zhou et al. (2024a) found it to be an effective pretraining strategy. Our SentenceTransformer FactFormer had an average of 21.4% less error across all of our data sets for standard training and average of 20.6% less error than baseline FactFormer when using transfer learning. Additionally, we see smaller improvement over baseline results when using Llama as our LLM, with some notable increases in error on the Shallow Water and Navier Stokes data sets.
@@ 5.2 Autoregressive Rollout
Second, we perform autoregressive rollout starting from our initial condition for 40 steps. This evaluates how well our model is able to temporally extrapolate beyond our training horizon. In this case, we use the models trained in section 5.1 with no additional fine-tuning. Accumulated rollout error is presented in table 1. We see that our SentenceTransformer FactFormer has lowest accumulated error for four of five data sets both for standard training and transfer learning, and outperforms baseline across all data sets. Our Llama FactFormer also outperforms baseline across all of our data sets except Shallow Water with standard training. For SentenceTransformer Factformer, we have an average reduction in accumulated error of 64.2% and 54.2% across all of our data set for standard training and transfer learning, respectively. For Llama FactFormer, we have an average reduction in accumulated error of 6.2% and 45.3% for standard and transfer learning, respectively. When removing Shallow Water for Llama Factformer with standard training, we have an average reduction in error of 63.1%.
We note that accumulated autoregressive rollout error for our Shallow Water benchmark is much higher for our transfer learning experiments than the standard training. This instability been noted before in Lorsung et al. (2024) when training errors are very low. In our case, training error for the Shallow Water data set is approximately an order of magnitude lower than other data sets. Error plots are given in figure 8 in Appendix C.
| Baseline | ST | Llama | Transfer | ST + Transfer | Llama + Transfer | |
| Combined | 1.36 0.03 | 0.78 0.04 | 0.90 0.02 | 1.30 0.08 | 0.80 0.09 | 0.86 0.04 |
| Burgers | 3.49 0.15 | 0.97 0.10 | 1.41 0.42 | 4.06 3.93 | 0.85 0.07 | 1.03 0.10 |
| Heat | 1.08 0.10 | 0.16 0.01 | 0.47 0.09 | 1.54 0.27 | 0.30 0.17 | 0.13 0.07 |
| Navier Stokes | 1.52 0.46 | 0.54 0.25 | 0.26 0.09 | 0.85 0.57 | 0.25 0.00 | 0.40 0.08 |
| Shallow Water | 0.12 0.09 | 0.05 0.04 | 0.35 0.36 | 52.20 50.63 | 51.22 60.16 | 65.70 21.50 |
5.3 Data Scaling
Using the same training setup as in section 5.1, we benchmark our multimodal models against baseline as we increase the amount of data for the Heat and Burgers data sets, given in figure 4, with scaling for the remaining datasets given in Appendix E. For the Heat and Burgers data sets, we see that our multimodal approach consistently outperforms baseline for our both large and small data settings, both with SentenceTransformer and Llama, and for both standard training and transfer learning. We add more data the relative improvement increases due to the plateau in our base models.

### v5

@@ Abstract
Solving Partial Differential Equations (PDEs) is ubiquitous in science and engineering. Computational complexity and difficulty in writing numerical solvers has motivated the development of data-driven machine learning techniques to generate solutions quickly. The recent rise in popularity of Large Language Models (LLMs) has enabled easy integration of text in multimodal machine learning models, allowing easy integration of additional system information such as boundary conditions and governing equations through text. In this work, we explore using pretrained LLMs to integrate various amounts of known system information into PDE learning. Using FactFormer as our testing backbone, we add a multimodal block to fuse numerical and textual information. We compare sentence-level embeddings, word-level embeddings, and a standard tokenizer across 2D Heat, Burgers, Navier-Stokes, and Shallow-Water data sets. These challenging benchmarks show that pretrained LLMs are able to utilize text descriptions of system information and enable accurate prediction using only initial conditions.
@@ 1 Introduction
Recently, machine learning methods have exploded in popularity to address these downsides in traditional numerical solvers. The primary aim is to reduce time-to-solution and bypass expensive calculations. Physics Informed Neural Networks (PINNs) (Raissi et al., 2019) have been widely successful in small-scale systems, but are notoriously difficult to train (Krishnapriyan et al., 2021). More recently, neural operators (Li et al., 2021; Lu et al., 2021; Li et al., 2023a) have improved upon PINNs, showing promise for larger scale, general purpose surrogate models. However, these neural operator models generally are purely data-driven, and do not use any additional, known, system information.
Additionally, owing to the success of large language models (LLMs) in many other domains, such as robotics (Kapoor et al., 2024; Bartsch & Farimani, 2024), and design (Kumar et al., 2023; Badagabettu et al., 2024; Jadhav & Farimani, 2024; Jadhav et al., 2024), some works have begun incorporating LLMs into the PDE surrogate model training pipeline. Universal Physics Solver (UPS) (Shen et al., 2024) incorporates pretrained LLMs, but uses limited text descriptions that do not fully utilize LLM capabilities. Unisolver (Zhou et al., 2024b) takes in a LaTeX description of the system as a prompt, but uses MLP encoders for additional system information such as boundary conditions, which does not fully explore the capabilities of the LLM. ICON-LM (Yang et al., 2024) uses longer text descriptions, but trains the LLM to make numerical predictions from input data and captions, adding additional complexity to the model architecture. Lastly, FMINT (Song et al., 2024) prompts an LLM to generate text descriptions, but only explores ODE systems. Both FMINT and ICON-LM use softmax-based attention which may not scale well to larger systems (Cao, 2021), as well.
Contributions: In this work, we explore differing levels of text description, fused with a mutlimodal block using pretrained LLMs. Word-level embeddings from Llama (AI@Meta, 2024) are popular for Natural Language Processing tasks and are compared against sentence-level embeddings from SentenceTransformer (Reimers & Gurevych, 2019), because global sentence information may be more useful in this case. These are compared against a tokenizer from Llema (Azerbayev et al., 2023) in order to evaluate how much the semantic knowledge of these pretrained LLMs affects downstream performance. To that end, we introduce a cross-attention based multimodal block that integrates textual information into FactFormer (Li et al., 2023b), given in figure 2. We benchmark this multimodal approach on the Heat, Burgers, Navier Stokes, and Shallow Water equations to provide a wide variety of physical behavior. Various boundary conditions, initial conditions, and operator coefficients are used, provide an additional challenge over existing benchmarks. We describe this system information with text, and incorporate this as conditioning information into FactFormer.
@@ 2 Related Work
Neural Solvers: Neural solvers have been developed largely to counter the drawbacks of traditional numerical methods. Physics Informed Neural Networks (PINNs) (Raissi et al., 2019) incorporate governing equations through a soft constraint on the loss function. PINNs have been shown to be effective in small-scale problems, but tend to be difficult to train (Wang et al., 2022; Rathore et al., 2024). Neural operators (Kovachki et al., 2023) were developed and show improvement over PINNs on a large variety of PDE learning tasks. Based on the universal operator approximation theorem, neural operators learn a functional that maps input functions to solution functions. Neural operators such as Fourier Neural Operator (FNO) (Li et al., 2021), based on kernel integral blocks in Fourier space, DeepONet(Lu et al., 2021), based on parameterizing the input functions and an embedding of the points at which the functions are evaluated, and OFormer (Li et al., 2023a), based on softmax-free attention. Despite very promising results, neural operators are usually purely data-driven and do not use any system information outside of solution fields.
Utilizing System Information While these neural operators perform well, they are often purely data driven. More recent works have incorporated various different aspects of the governing systems. Takamoto et al. (2023a) incorporates operator coefficients through the CAPE module. While coefficients are important in determining system behavior, system parameters such as boundary conditions, forcing terms, and geometry can play an equally important role, and cannot be easily incorporated through CAPE. Further work incorporates governing equations into models. Lorsung et al. (2024) developed the PITT framework that uses a transformer-based architecture to process system information as text. Additionally, PROSE (Liu et al., 2023) poses a multi-objective task to both make a prediction and complete partially correct governing equation, and follow up work in PROSE-FD (Liu et al., 2024) is similar to PITT in incorporating symbolic representations as input. These frameworks perform well, but rely on notational consistency between samples, and do not offer an easy integration of different geometries. Lastly, Hao et al. (2023) introduced novel Heterogeneous Normalized Attention and Geometric Gating mechanisms for flexible GNOT model. GNOT is able to incorporate different system information, such as coefficients and geometry. This flexible architecture can incorporate many different system parameters, but requires additional implementation details for each additional modality, and may not be able to capture qualitative aspects of systems, such as flow regime.
@@ 3 Data Generation
We generate data from the Heat, Burgers, and Incompressible Navier Stokes equations, which are popular benchmarks for fluid dynamics surrogate models. Additionally, we benchmark on the Shallow Water equations from PDEBench (Takamoto et al., 2023b). Data for the Heat and Burgers equations are generated using Py-PDE (Zwicker, 2020) due to ease of simulating different boundary conditions, and data is generated for Navier Stokes using code from Fourier Neural Operator (Li et al., 2021). A diverse data set is generated in order to create a more challenging setup than existing benchmarks, that often have the same operator coefficients and boundary conditions for all samples, only varying the initial conditions. In our case, we use different initial conditions, operator coefficients, and boundary conditions. While existing data sets offer distinct challenges with regards to governing equations, they do not offer a lot of data diversity with regard to system parameters, which LLMs are well-suited to handle. PDEBench, for example, has multiple 2D data sets that represent a variety of physical processes. However, only eight different operator coefficient combinations are used for the 2D Compressible Navier Stokes equations, all with the same boundary conditions. In order to fully utilize the capabilities of pretrained LLMs, as well as present a more challenging benchmark, we vary boundary conditions and operator coefficients for the Heat and Burgers equations in 2D, and vary viscosity and forcing term amplitude for the Incompressible Navier-Stokes equations in 2D.
@@ 3.1 Heat Equation
We generate data on a simulation cell given by: on a 64x64 grid with boundary conditions sampled from . The values for the Neumann and Dirichlet boundary conditions are sampled uniformly: . All four walls have the same boundary condition type and value for a given simulation. Initial conditions are given by Lastly, our diffusion coefficient was sampled randomly according to . We simulated each trajectory for 1 second on a 64x64 grid that was evenly downsampled to 100 frames, for 101 total snapshots of data including initial conditions. These distribution of diffusion coefficients was chosen so the diffusive dynamics behaved on approximately the same scale as the advection dynamics of Burgers Equation given below.
@@ 3.2 Burgers’ Equation
Burgers equation models shock formation in fluid waves, given below in equation 2, where we are predicting the height, at each time step.
@@ 3.3 Incompressible Navier-Stokes Equations
The Navier-Stokes data generated here follows the setup from (Lorsung et al., 2024) with different viscosities and forcing term amplitudes. Our simulation cell is given by: with periodic boundary conditions. Our viscosity is sampled according to
, and our forcing term amplitude is sampled according to . The initial condition is given by a Gaussian random field and out simulation is run for 10 seconds. Data generation is done on a grid that is evenly downsampled to a 64x64 grid for this work, and 100 evenly sampled timesteps with initial condition, for 101 total snapshots of data.
@@ 3.4 Shallow-Water Equations
Lastly, we take the Shallow-Water data set from PDEBench (Takamoto et al., 2023b). In this case, we are predicting the water height at each time , where is x-velocity, is y-velocity, is gravity, and is the bathymetry.
@@ 4.1 System Descriptions
All of our basic system descriptions come from Wikipedia entries for their respective equations (Wikipedia, 2024a, b, c, d). Next, boundary condition information is added, in this case Neumann boundary conditions:
@@ 4.2 Multimodal PDE Learning
The backbone of our multimodal surrogate model is the FactFormer (Li et al., 2023b) and our framework is given in figure 1. FactFormer was chosen because it provides a fast and accurate benchmark model. We add our system description as conditioning information both before and after factorized attention blocks. FactFormer is a neural operator that learns a functional that maps from input function space to solution function space as , with parameters . In our case, we are specifically learning the functional conditioned on our system information . For a given input function evaluated at points , we are learning the the operator given in equation 5. That is, our network learns to make predictions for our solution function also evaluated at points .
@@
4.3 Text Processing
This multimodal block is agnostic to both the data-driven backbone as well as the LLM. We compare Llama 3.1 8B (AI@Meta, 2024), and all-MiniLM-L6-v2 from the Sentence-Transformer (ST) package (Reimers & Gurevych, 2019) as our pretrained LLMs, as well as the tokenizer from Llema (Azerbayev et al., 2023). Llama 3.1 8B was used because of its good performance across a wide variety of benchmarks, as well as its small size allowing us to use only a single GPU to generate sentence embeddings. The word embeddings are averaged for each sentence to provide a single embedding that can be used in our projection layer. all-mpnet-base-v2 was used because it is designed to generate embeddings from sentences that are useful for tasks like sentence similarity. This allows downstream tasks to utilize embeddings that utilize the entire sentence, similar to human reasoning at the sentence level. Lastly, our tokenizer-only variant is used to provide insight in how much semantic knowledge our models are able to utilize from our pretrained LLMs, because a tokenizer alone has none. In all of our experiments, the LLM is frozen and a projection head is trained. This significantly improves training time by allowing us to generate the sentence embeddings before training, and avoiding expensive gradient computations for our LLMs.
@@ 5 Results
We benchmark our multimodal model against its baseline variant on a number of challenging tasks. Our data vary the distribution of initial conditions, operator coefficients, and boundary conditions, which provides a much more challenging setting than many existing benchmarks. In each experiment, the combined data set is the Heat, Burgers, and Navier-Stokes data sets, where Shallow Water is used solely for fine-tuning, with a 80-10-10 train-validation-test split. In our transfer learning experiments, we first train on this combined data set, then finetune on each individual data set. Transfer learning was done because Zhou et al. (2024a) found it to be an effective pretraining strategy. 1,000 samples per equation are used for each data set. The effect of text information is evaluated on next-step prediction, autoregressive rollout, and fixed-future prediction tasks. We use only the initial condition as data input in all of our experiments. Using only a single step of input here presents the additional challenge that it is difficult to infer boundary conditions and impossible to infer operator coefficients from only a single frame of data. Relative Error (Li et al., 2021) is used for both training and next-step evaluation in all of our experiments. Autoregressive rollout error is reported as Mean Squared Error. Experiments use various combinations of boundary condition (B), coefficient (C) and qualitative (Q) information. Each reported result is the mean and standard deviation across three random seeds. Results for our combined data sets, as well as tabular values for training are given in appendix C. Training details are given in appendix B.
@@ 5.1 Next-Step Prediction
First, we evaluate next-step predictive accuracy, where we take one frame of data and use it to predict the next across our entire temporal window. We see in figure 3 that our multimodal FactFormer significantly outperforms baseline FactFormer for each of our multimodal approaches. In this case, SentenceTransformer either has best performance, or is close to best performance across all of our data sets with both standard training and transfer learning. Both our tokenizer and Llama variants offer improvement over baseline as well. Because our Shallow Water data set does not have varying coefficients, boundary conditions, or qualitative behavior, we see comparable performance between baseline and multimodal models. This is corroborated in our ablation study in section 5.3. When excluding the Shallow-Water results, our SentenceTransformer variant has an average reduction in error of 39.0% and 47.6% for standard training and transfer learning, respectively, when compared to our baseline method.
@@ 5.2 Autoregressive Rollout
Second, we perform autoregressive rollout starting from our initial condition for 40 steps. This evaluates how well our model is able to temporally extrapolate beyond our training horizon. In this case, we use the models trained in section 5.1 with no additional fine-tuning. Accumulated rollout error is presented in figure 4. Our autoregressive rollout error results follow closely from the next-step prediction results. We see that our SentenceTransformer variant has lowest error for our Heat, Burgers, and Navier Stokes data sets for both standard training and transfer learning. On our Shallow Water data sets, we note comparable performance for all models for the reasons noted in section 5.1. Similarly, when excluding the Shallow Water results, the SentenceTransformer variant has an average reduction in accumulated error of 67.4% and 81.3% for standard training and transfer learning, respectively.
We note that accumulated autoregressive rollout error for our Shallow Water benchmark is much higher for our transfer learning experiments than the standard training. This instability been noted before in (Lorsung et al., 2024) when training errors are very low. In our case, training error for the Shallow Water data set is approximately an order of magnitude lower than other data sets.
5.3 Fixed-Future
In this case, we use the initial condition to predict the final simulation state for each data set. Results are given in figure 5. We see clear improvement over baseline in the Heat, Burgers, and Navier-Stokes data sets for each of our multimodal variants, with marginal improvement over Shallow Water for the same reasons given in section 5.1. Interestingly, we see that our Llama variant has greatest improvement over baseline, in contrast to the next-step and autoregressive rollout experiments. Our Llama variant has an average reduction in error of 50.2% and 57.1% when excluding Shallow Water for standard training and transfer learning, respectively.